Grade 11IIT JEE Entrance Exam

The phase difference between two point separated by 0.8m in a wave of frequency 100 Hz is π\ 2 the wave velocity is

Sudhanva G V
7 Years agoGrade 11

1 Answer

TARUN SHARMA
7 Years ago
 phi=pi/2, n=100hz.
 
as we know lamda = velocity/ frequency.
 
phi = 2pi x/lamda
   pi / 2 = 2 pi 0.8/lamda
from above solving lamda = 4 * 0.8 
  lamda = 3.2
 
thus lamda = v / n
  v(velocity) = n * lamda 
 
   v = 100*3.2
 v= 320 m/s